OpenStudy (anonymous):

solve the integral (14x+1)/(4x^2-13x-12) .. please help guys .. no one is answering me lately =[ and tomorrow i have a test !!!!

OpenStudy (amistre64):

is the top the derivative of the bottom?

OpenStudy (anonymous):

nope ... because you have 13x downstairs and on top 1

OpenStudy (amistre64):

good, then most likely we have to do a decomp whats the fators of the bottom?

OpenStudy (amistre64):

factors even

OpenStudy (anonymous):

i didin't manage to find real roots from the quadratic equation ...

OpenStudy (amistre64):

let me chk that then 4x^2-13x-12 13^2-4(4)(-12) = 196+ some value it doesnt go complex

OpenStudy (anonymous):

oh yes you are right !!!! i did a mistake with the minus sign !!! any way can you just tell me how to factor when you find the roots ? =]

OpenStudy (turingtest):

\[\frac{14x+1}{4x^2-13x-12}=\frac A{4x+3}+\frac B{x-4}\]

OpenStudy (amistre64):

16 and 3 (x-16/4)(x+3/4) (x-4)(4x+3)

OpenStudy (amistre64):

youhave a test, how is just telling you going to give you the means to solve it? that comes from experience .... hands on stuff

OpenStudy (amistre64):

turing gave the split up; now we have to solve for A and B

OpenStudy (amistre64):

we do that by trying to recombine the fractions using the unknowns and comparing it with the known numberator

OpenStudy (anonymous):

yes ... i see ... thank you very much !!

OpenStudy (amistre64):

14x + 1 = A(x-4) + B(4x+3) the simple means here is to zero one of the terms out; when x=4 we get A(0) sooo 14(4) + 1 = B(4(4)+3) B = 57/19 if i see that right :)

OpenStudy (amistre64):

B cancels out when x = -4/3

OpenStudy (anonymous):

yes !! thank you i can go on from here peacfully =] .. i will be happy for help with my other questions from the last 2 hours that no one replied yet ;]

OpenStudy (amistre64):

if you repost it we will be able to find it easier, with any luck :)